OpenStudy (anonymous):

PLEASE HELP!!! A pair of skis cost $80 more than a pair of boots, together they sell for $294.25 after a 7% tax rate. what is the equation and the cost of the skis.

OpenStudy (anonymous):

jthuga

OpenStudy (mathteacher1729):

Let's break this down piece by piece: "A pair of skis cost $80 more than a pair of boots." Let B = Price of Boots Let S = Price of Skis \[S = 80+B\] "Together they sell for $294.25" \[B + S = 294.25\] But S = 80 + B, so re-write: \[B + (B+80)=294.25\] Ok, that would be our setup IF... there were no Tax!! :( So we know that (B + (B+80)) has been marked up by 7%, or in other words, multiplied by 1.07. The final equation is then this: \[1.07(B + (B+80))=294.25\] It says "Take the price of the boots, B, plus the price of the skis (B =80) and take that whole sum, and raise it by 7% (or multiply by 1.07). All of that is equal to $294.25." So now we have to solve... From here it's an algebra problem where you solve for B. Remember though, when you find B you're NOT DONE! You still have to add $80 to it to find the price of the skis. Hope this helps!